# What Language Models Know But Don’t Say: Non-Generative Prior Extraction for Generalization

### v1

@@ Abstract
In domains like medicine and finance, large-scale labeled data is costly and often unavailable, leading to models trained on small datasets that struggle to generalize to real-world populations. Large language models contain extensive knowledge from years of research across these domains. We propose LoID (Logit-Informed Distributions), a deterministic method for extracting informative prior distributions for Bayesian logistic regression by directly accessing their token-level predictions. Rather than relying on generated text, we probe the model’s confidence in opposing semantic directions (positive vs. negative impact) through carefully constructed sentences. By measuring how consistently the LLM favors one direction across diverse phrasings, we extract the strength and reliability of the model’s belief about each feature’s influence. We evaluate LoID on ten real-world tabular datasets under synthetic out-of-distribution (OOD) settings characterized by covariate shift, where the training data represents only a subset of the population. We compare our approach against (1) standard uninformative priors, (2) AutoElicit, a recent method that prompts LLMs to generate expert-style priors via text completions, (3) LLMProcesses, a method that uses LLMs to generate numerical predictions through in-context learning by formatting tabular data as natural language prompts and (4) an oracle-style upper bound derived from fitting logistic regression on the full dataset. We assess performance using Area Under the Curve (AUC). Across datasets, LoID significantly improves performance over logistic regression trained on OOD data, recovering up to 59% of the performance gap relative to the oracle model. LoID outperforms AutoElicit and LLMProcessesc on 8 out of 10 datasets, while providing a reproducible and computationally efficient mechanism for integrating LLM knowledge into Bayesian inference.
What Language Models Know But Don’t Say:
Non-Generative Prior Extraction for Generalization
Sara Rezaeimanesh Michigan State University rezaeima@msu.edu          Mohammad M. Ghassemi Michigan State Unviersity ghassem3@msu.edu
@@ 1 Introduction
Machine learning models used in critical domains like medicine and finance often lack access to large, well-balanced, and comprehensively labeled datasets. Clinical datasets may cover only a single hospital or demographic, while credit-risk tables can be skewed toward historical lending patterns. For example, Purushotham et al. (2017) benchmarked several deep learning models on clinical prediction tasks—including mortality and length of stay—using the MIMIC-III intensive care electronic health record dataset from a single hospital site. As a consequence, models trained on small, non-representative samples frequently fail to generalize when confronted with real-world populations that differ from their training distribution Brigato and Iocchi (2020); Li et al. (2024); Althnian et al. (2021). Bayesian methods offer a principled way to encode prior knowledge and mitigate such brittleness, but manually eliciting high-quality priors at scale remains a longstanding challenge O’Hagan et al. (2006).
@@
Parallel to this, out-of-distribution (OOD) generalization has gained increasing attention, as real-world deployment often involves data that differs from the development environment. Meta-learning approaches like TabPFN Hollmann et al. (2023) and Drift-Resilient TabPFN Helli et al. (2024) address OOD robustness through synthetic task generation and causal modeling. Others have proposed logit-based calibration strategies such as UE-NL Huang and Qiao (2023), or LLM-guided data augmentation methods like Summary Boosting Manikandan et al. (2023) and LLMOverTab Isomura et al. (2025) to cover underrepresented regions of the feature space. However, despite this momentum, few works have systematically explored how to extract priors from the internal structure of LLMs (e.g. hidden states, logits, etc.) for tabular tasks under distribution shift.
We address this gap with LoID (Logit-Informed Distributions), a deterministic framework that elicits informative priors for Bayesian logistic regression directly from an LLM’s token-level logits. Instead of relying on generated completions or examples, LoID probes the LLM with semantically paired sentences highlighting positive and negative feature-target relationships. The logit difference between the relational tokens quantifies directional belief, and the variance across paraphrased prompts yields an uncertainty estimate—together forming Normal priors over model coefficients. These priors integrate into Bayesian workflows, without any fine-tuning or in-context learning.
We evaluate LoID on ten publicly available tabular datasets that are widely used benchmarks and have appeared in numerous prior studies, ensuring comparability with existing methods. We compare LoID against uninformative Bayesian inference and LLMProcesses. To provide an upper bound, we additionally train and evaluate a logistic regression model on the entire dataset, offering insight into performance when full population data is available. Across extensive out-of-distribution splits, LoID outperforms existing approaches in predictive accuracy (AUC). By deterministically extracting domain knowledge encoded in large language models, LoID enables robust and reproducible inference.
@@
  • Logit-based deterministic prior elicitation. LoID is the first method to extract Bayesian priors directly from the token-level logits of large language models, eliminating the variability associated with text generation and prompt sampling.
  • Baseline comparison across ten OOD datasets. We evaluate LoID on diverse OOD tabular datasets from healthcare, finance, and social science. Our experiments demonstrate that LoID often outperforms existing approaches across most settings.
@@ 2.1 LLMs for Tabular Data Modeling
Beyond standalone applications, recent research has focused on leveraging LLMs to enhance classical statistical models. For instance, Abhyankar et al. (2025) introduced LLM-FE, a framework that combines evolutionary search with the reasoning capabilities of LLMs to iteratively generate and refine feature transformations for tabular data. Similarly, LLM-Lasso Zhang et al. (2025) is a domain-informed approach to feature selection and regularization by leveraging large language models to guide the selection process. Capstick et al. (2025) proposed AutoElicit, a method that interacts with LLMs to extract priors for Gaussian-based probabilistic models; however, this approach primarily focuses on Gaussian priors and lacks flexibility in handling other priors that may be more suitable for different data characteristics. In a different vein, Zhong et al. (2024) presented LLMCorr, a training-free framework where LLMs act as post-hoc correctors to propose corrections for the predictions of arbitrary machine learning models. Liu et al. (2023) investigated the reasoning capabilities of LLMs over tabular data and proposed combining symbolic and textual reasoning to improve performance on table-based question answering tasks.
@@ 2.2 Out-of-Distribution Generalization
Our work, LoID, complements and extends this growing body of research by proposing the first deterministic logit-based method to elicit Bayesian priors from LLMs for structured tabular tasks. Our method consistently improves OOD generalization across ten tabular datasets, offering a lightweight and almost deterministic integration of LLM knowledge into classical statistical pipelines.
@@ 3 Method
We propose LoID (Logit-Informed Distribution), a deterministic framework for extracting priors for tabular dataset features from LLMs using token-level logits, rather than relying on generated samples. Our goal is to derive informative priors for Bayesian logistic regression that generalize robustly in low-data settings.
@@ 3.1 Problem Setting
Let $\mathcal{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote a training dataset, where each $x_{i}\in\mathbb{R}^{d}$ is a vector of $d$ features and $y_{i}\in\{0,1,\ldots,m\}$ is the corresponding label. In Bayesian logistic regression, we model the likelihood as
@@
where $\sigma(\cdot)$ denotes the sigmoid function and $\boldsymbol{\beta}\in\mathbb{R}^{d}$ represents the regression coefficients. Bayesian inference is performed by specifying a prior distribution $p(\boldsymbol{\beta})$ and computing the posterior $p(\boldsymbol{\beta}\mid\mathcal{D}_{\text{train}}).$ In data-scarce regimes, informative priors can substantially improve generalization. LoID constructs such priors by leveraging domain knowledge encoded in large language models, without access to the training data.
@@ 3.2 Querying LLM Beliefs via Logit Differences
For each feature $x_{j}$, we construct paired prompts that express opposing semantic claims about its effect on the target variable. Each prompt is carefully designed such that all relevant context appears before a sentiment-bearing token (e.g., “positive” or “negative”), ensuring that the token’s probability considers the entire context for the task, including the name of the feature and the prediction task (see examples in A.2.1). For example, given a prompt prefix like:
@@
we use the probabilities from the LLM’s softmax over its vocabulary to extract the probability of the next token being “positive” or “negative”. We define:
$\mathcal{P}_{j}^{+}$ $=P(\texttt{``positive''}\mid\text{prefix})$
$\mathcal{P}_{j}^{-}$ $=P(\texttt{``negative''}\mid\text{prefix})$
where $f_{j}$ is a natural language description of feature $x_{j}$ (e.g., “age of the patient”) and $t$ describes the prediction task (e.g., “disease risk”).
We define the logit preference score for a single prompt $p_{j}$ as:
$\text{logit}(p_{j})=\log\left(\frac{p_{j}}{1-p_{j}}\right)$
Where
$p_{j}=\frac{P_{j}^{+}}{P_{j}^{+}+P_{j}^{-}}$
The logit preference score quantifies the extent to which the model prefers the token “positive” over “negative” in a given context. While this difference may also reflect other patterns learned during training, our hypothesis is that by constructing sentences that do not grammatically or syntactically favor either token, we reduce the influence of such confounding factors. As a result, the observed preference is driven primarily by the semantic relationship between the feature and the target concept. Finally, applying the logit transformation maps probabilities from the bounded interval (0, 1) to the unbounded real line, making the resulting scores directly appliable as coefficients in logistic regression models.
The mean prefrence score $\mu_{j}$ for feature $x_{j}$ is given by averaging this score across a set of paraphrased prompts T{T}:
$\mu_{j}=\frac{1}{|T|}\sum_{t\in T}\text{logit}_{t}(p_{j})$
This value reflects the LLM’s normalized preference for a positive versus negative association with the feature across diverse prompt formulations. Figure 1 demonstrates how LoID extracts informative priors from model logits.
@@ 3.3 Constructing Priors from Logits
We model each coefficient $\beta_{j}$ with a Gaussian prior: $\beta_{j}\sim\mathcal{N}(\mu_{j},\sigma_{j}^{2})$. The mean $\mu_{j}$ is the average logit preference score across paraphrases, which reflects the direction and strength of influence the LLM assigns to feature $x_{j}$. To capture uncertainty in the LLM’s belief, we define the standard deviation $\sigma_{j}$ as a scaled function of variability in the individual logit preference scores:
$\sigma_{j}^{2}=\alpha+\gamma\cdot\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
Here, $logit_{t}(p_{j})$ denotes the logit preference score for prompt $t$ and feature $j$. $\alpha>0$ and $\gamma>0$ are fixed scaling hyperparameters that can be tuned to optimize LoID’s performance on a certain task. We perform a comprehensive hyperparameter tuning in A.2.4 to find the hypermarameters that achieve the highest cumulative AUC. This design assigns higher variance to features for which the LLM expresses inconsistent preferences across paraphrases, thereby reflecting greater uncertainty about the feature’s influence in the model.
@@ 3.4 Bayesian Inference with LoID Priors
With the LoID priors in place, posterior inference is performed using the No-U-Turn Sampler (NUTS), a variant of Hamiltonian Monte Carlo (HMC). HMC augments the parameter space with auxiliary momentum variables and explores the posterior by simulating Hamiltonian dynamics, enabling efficient, gradient-informed proposals that avoid the random-walk behavior of traditional MCMC methods.
Specifically, HMC introduces a momentum variable $\mathbf{r}$ for the regression coefficients $\boldsymbol{\beta}$ and defines the Hamiltonian
$\mathcal{H}(\boldsymbol{\beta},\mathbf{r})=-\log p(\boldsymbol{\beta}\mid\mathcal{D}_{\text{train}})+\frac{1}{2}\mathbf{r}^{\top}\mathbf{r}.$
Proposals are generated by numerically simulating the resulting Hamiltonian dynamics using a leapfrog integrator, which leverages gradients of the log-posterior to propose distant yet highly probable states.
NUTS eliminates the need to manually tune the trajectory length by adaptively extending the simulation until a U-turn condition is detected, ensuring efficient exploration while maintaining detailed balance. This results in faster mixing and improved effective sample sizes, particularly in high-dimensional parameter spaces.
We run multiple NUTS chains with warm-up steps for step-size and mass-matrix adaptation, and discard burn-in samples before collecting posterior draws. All reported results are computed using posterior expectations estimated from these samples.
3.5 Summary
LoID transforms the implicit domain knowledge encoded in large language models into explicit, feature-level priors suitable for probabilistic modeling. Unlike prior approaches that rely on generated samples, in-context learning, or fine-tuning, LoID achieves this deterministically by directly probing token-level logits. This design enables reproducibility and principled integration into Bayesian models.
@@ 3.6 Evaluation
To assess the quality of the resulting posteriors, we compute the AUC score on the entire dataset using models trained solely on the OOD training split. We then compare this AUC to that of a logistic regression trained on the entire dataset. This comparison quantifies how much of the predictive performance gap, caused by training on an OOD distribution, is recovered by our method.
@@ 4.1 Datasets
We evaluate and compare LoID’s performance on ten publicly available tabular datasets for binary and multi-class classification: The Heart Disease (Cleveland) dataset Janosi et al. (1989) predicts the presence of heart disease based on clinical and demographic attributes such as cholesterol level, and blood pressure. The Indian Liver Patient Dataset (ILPD) Ramana and Venkateswarlu (2022) classifies whether a patient has liver disease using demographic information and blood test measurements. The Blood Transfusion Service Center dataset Yeh (2008a) predicts whether a donor will donate blood again based on historical donation behavior. The Room Occupancy Detection dataset Candanedo (2016b) determines whether a room is occupied using time-series environmental sensor readings.
The Diabetes dataset Islam et al. (2020) predicts the likelihood of diabetes onset based on self-reported symptoms. The Pima Indians Diabetes dataset Smith et al. (1988) predicts the presence of diabetes using diagnostic measurements such as glucose concentration and BMI. The Give Me Some Credit (Good Credit Score) dataset Fusion and Cukierski (2011) assesses an individual’s creditworthiness based on income, debt ratio, and delinquency history. The Bank Marketing dataset Moro et al. (2014) predicts whether a client will subscribe to a term deposit based on demographic attributes and marketing interactions.
The Adult Income (Census Income) dataset Becker and Kohavi (1996b) predicts whether an individual’s annual income exceeds \$50K using demographic and educational features. Finally, the Jungle Chess (Jungle) dataset Shapiro (1983) predicts game outcomes based on structured representations of board states, providing a symbolic decision-making task.
For each dataset, we construct synthetic OOD splits as described in Section A.4.2. To verify that the resulting training sets do not reflect the true data distribution, a logitic regression model is trained separately on the out-of-distribution and in-distribution training splits and evaluated on the entire dataset. As shown in Table 1, the model trained on the OOD train set consistently underperforms compared to the model trained on the full dataset. This confirms that generalization is impaired and the OOD train set does not present the full distribution of the dataset as well as the in distribution train set.
@@
Across all datasets, we apply the same preprocessing pipeline: categorical features are one-hot encoded, and missing values are imputed with zeros. For datasets with an excessive number of features—namely Give Me Some Credit, and Adult Income—we perform feature selection using GPT-4o by providing feature descriptions and prompting it to return a list of the 15 most relevant features for the prediction task. The same reduced feature set is used consistently across all experimental conditions to ensure comparability.
Dataset AUC (IID) AUC (OOD)
Heart Disease 0.93 0.86
Liver Disease 0.77 0.70
Blood Donation 0.75 0.58
Room Occupancy 1.00 0.98
Diabetes 0.84 0.81
Income 0.91 0.78
Credit 0.79 0.71
Jungle 0.81 0.71
Pima Diabetes 0.84 0.81
Bank Marketing 0.89 0.51
Table 1: AUC comparison of logistic regression on IID vs. OOD test splits. Lower performance on OOD indicates distribution shift.
@@
We designate the Heart Disease, Blood Donation, and Diabetes datasets as development datasets for evaluating prompt structures and hyperparameter tuning. The remaining five datasets are reserved for testing the generalization performance of the final LoID method. See (A.2.4) for details on the hyperparameter search process and value ranges explored.
Dataset OOD-LR AutoElicit (baseline) LLMProc (baseline) LoID (our method) Cap (upper bound) Gap Closed%
Heart Disease 0.87 0.92 0.86 0.90 0.93 +50
Liver Disease 0.70 0.36 0.60 0.69 0.77 -19
Blood Donation 0.58 0.71 0.63 0.67 0.75 +52
Room Occupancy 0.98 0.97 0.80 1.00 1.00 +78
Diabetes 0.81 0.83 0.70 0.84 0.84 +99
Good Credit Score 0.71 0.71 0.30 0.76 0.79 +60
Income 0.78 0.85 0.51 0.87 0.91 +57
Pima Diabetes 0.82 0.83 0.74 0.84 0.84 +88
Bank Marketing 0.51 0.51 0.51 0.72 0.89 +56
Jungle Chess 0.71 0.78 0.62 0.79 0.81 +70
Table 2: Comparison of AUC scores across datasets. “OOD-LR” reports the performance of logistic regression trained on an OOD training set. LoID results are obtained using priors from Gemma-2-27B, with $\alpha=0.2$, $\gamma=2.0$ and $N_{sent}=10$. Among Uniformative, AutoElicit, LLMProcesses, and LoID, the highest AUC for each dataset is shown in bold. The Cap condition serves as an oracle-style upper bound. All values are rounded to two decimal places. In some cases, LoID and Cap appear to have similar performance; however, the gap is not fully closed. This is due to rounding. For example, on Diabetes, LoID achieves 0.839 while Cap achieves 0.842.

### v2

# What LLMs Know but Don’t Say: Non-generative Prior Extraction for Generalization

@@ Abstract
In domains like medicine and finance, large-scale labeled data is costly and often unavailable, leading to models trained on small datasets that struggle to generalize to real-world populations. Large language models contain extensive knowledge from years of research across these domains. We propose LoID (Logit-Informed Distributions), a deterministic method for extracting informative prior distributions for Bayesian logistic and linear regression by directly accessing their token-level predictions. Rather than relying on generated text, we probe the model’s confidence in opposing semantic directions (positive vs. negative impact) through carefully constructed sentences. By measuring how consistently the LLM favors one direction across diverse phrasings, we extract the strength and reliability of the model’s belief about each feature’s influence. We evaluate LoID on 15 real-world tabular datasets under synthetic out-of-distribution (OOD) settings characterized by covariate shift. We compare our approach against (1) standard uninformative priors, (2) AutoElicit, a method that prompts LLMs to generate expert-style priors, (3) LLMProcesses, a method that prompts LLMs to generate predictions and (4) an oracle-style upper bound derived from fitting logistic/linear regression on an in-distribution (IID) train set. We assess performance using Area Under the Curve (AUC) and Mean Squared Error (MSE). Across datasets, LoID significantly improves performance over logistic and linear regression models trained on OOD data, recovering up to 50% of the performance gap relative to the oracle model.
@@ 1 Introduction
Machine learning models used in critical domains like medicine and finance often lack access to large, well-balanced, and comprehensively labeled datasets. Clinical datasets may cover only a single hospital or demographic, while credit-risk tables can be skewed toward historical lending patterns. For example, Purushotham et al. (2017) benchmarked several deep learning models on clinical prediction tasks—including mortality and length of stay—using the MIMIC-III intensive care electronic health record dataset from a single hospital site. As a consequence, models trained on small, non-representative samples frequently fail to generalize when confronted with real-world populations that differ from their training distribution Brigato and Iocchi (2020); Li et al. (2024); Althnian et al. (2021). Bayesian methods offer a principled way to encode prior knowledge and mitigate such brittleness, but manually eliciting high-quality priors at scale remains a longstanding challenge Mikkola et al. (2023).
@@
While LLMs offer a promising source of priors for downstream modeling, their practical benefit depends on how well these priors generalize under distribution shift. OOD generalization has therefore received increasing attention, as real-world deployment often involves data that differs from the development environment. Meta-learning approaches like TabPFN Hollmann et al. (2023) and Drift-Resilient TabPFN Helli et al. (2024) address OOD robustness through synthetic task generation and causal modeling. Others have proposed logit-based calibration strategies such as UE-NL Huang and Qiao (2023), or LLM-guided data augmentation methods like Summary Boosting Manikandan et al. (2023) and LLMOverTab Isomura et al. (2025) to cover underrepresented regions of the feature space. However, despite this momentum, few works have systematically explored how to extract priors from the internal structure of LLMs (e.g. hidden states, logits, etc.) for tabular tasks under distribution shift.
We address this gap with LoID (Logit-Informed Distributions), a deterministic framework that elicits informative priors for Bayesian logistic and linear regression directly from an LLM’s token-level logits. Instead of relying on generated completions or examples, LoID probes the LLM with semantically paired sentences highlighting positive and negative feature-target relationships. The logit difference between the relational tokens quantifies directional belief, and the variance across paraphrased prompts yields an uncertainty estimate—together forming Normal priors over model coefficients. These priors integrate into Bayesian workflows, without any fine-tuning or in-context learning.
We evaluate LoID on 14 publicly available tabular datasets that are widely used benchmarks and have been studied in prior work, along with one non-public medical dataset to assess performance on data that does not exist in pre-training corpora. We compare LoID against uninformative Bayesian inference, LLMProcesses, and AutoElicit. To provide an upper bound, we additionally train logistic and linear regression models on IID training sets, offering insight into performance when the true data distribution is available.
Our experiments show that LoID improves OOD generalization across 11/15 tabular datasets, recovering 50% of the lost performance on average.
@@
  • Logit-based deterministic prior elicitation. LoID is the first method to extract Bayesian priors directly from the token-level logits of LLMs, eliminating the variability associated with text generation and prompt sampling.
  • Baseline comparison across 15 OOD datasets. We evaluate LoID on diverse OOD tabular datasets from healthcare, finance, and social science. Our experiments demonstrate that LoID often outperforms existing approaches across most settings.
@@ 2.1 LLMs for tabular data modeling
Beyond standalone applications, recent research has focused on leveraging LLMs to enhance classical statistical models. For instance, Abhyankar et al. (2025) introduced LLM-FE, a framework that combines evolutionary search with the reasoning capabilities of LLMs to iteratively generate and refine feature transformations for tabular data. Similarly, LLM-Lasso Zhang et al. (2025) is a domain-informed approach to feature selection and regularization by leveraging large language models to guide the selection process. Capstick et al. (2025) proposed AutoElicit, a method that interacts with LLMs to extract priors for Gaussian-based probabilistic models. In a different vein, Zhong et al. (2024) presented LLMCorr, a training-free framework where LLMs act as post-hoc correctors to propose corrections for the predictions of arbitrary machine learning models. Liu et al. (2023) investigated the reasoning capabilities of LLMs over tabular data and proposed combining symbolic and textual reasoning to improve performance on table-based question answering tasks.
@@ 3 Method
We propose LoID (Logit-Informed Distribution), a deterministic framework for extracting priors for tabular dataset features from LLMs using token-level logits, rather than relying on generated samples. Our goal is to derive informative priors for Bayesian logistic and linear regression that generalize robustly in low-data settings.
@@ 3.1 Problem setting
Let $\mathcal{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote a training dataset, where each $x_{i}\in\mathbb{R}^{d}$ is a vector of $d$ features and $y_{i}\in\{0,1,\ldots,m\}$ is the corresponding label. In Bayesian logistic regression, we model the likelihood as:
@@
where $\sigma(\cdot)$ denotes the sigmoid function and $\boldsymbol{\beta}\in\mathbb{R}^{d}$ represents the regression coefficients. Bayesian inference is performed by specifying a prior distribution $p(\boldsymbol{\beta})$ and computing the posterior $p(\boldsymbol{\beta}\mid\mathcal{D}_{\text{train}}).$ In data-scarce settings, informative priors can substantially improve generalization. LoID constructs such priors by leveraging domain knowledge encoded in LLMs, without access to the training data.
@@ 3.2 Querying LLM beliefs via logit differences
For each feature $f_{j}$, we construct paired prompts that express opposing semantic claims about its effect on the target variable. Each prompt is carefully designed such that all relevant context appears before a sentiment-bearing token (e.g., “positive” or “negative”), ensuring that the token’s probability considers the entire context for the task, including the name of the feature and the prediction task, $t$ (see examples in A.2.1). For example, given a prompt prefix like:
@@
We use the probabilities from the LLM’s softmax over its vocabulary to extract the probability of the next token being positive or negative. For the $i$th prompt prefix we define:
$\mathcal{P}_{i}^{+}=P(\texttt{<positive\_token>}\mid\text{prefix}),\quad\mathcal{P}_{i}^{-}=P(\texttt{<negative>\_token}\mid\text{prefix})$
We then define the logit preference score as:
$\text{logit}_{j}(p_{i})=\log\left(\frac{p_{i}}{1-p_{i}}\right),\quad p_{i}=\frac{P_{i}^{+}}{P_{i}^{+}+P_{i}^{-}}$
The logit preference score quantifies the extent to which the model prefers the positive token over the negative token in a given context. While this difference may also reflect other patterns learned during training, our hypothesis is that by constructing sentences that do not grammatically or syntactically favor either token, we reduce the influence of such confounding factors. Finally, applying the logit transformation maps probabilities from the bounded interval (0, 1) to the unbounded real line, making the resulting scores directly appliable as coefficients in logistic regression models.
@@ 3.3 Constructing priors from logits
We model each coefficient $\beta_{j}$ with a Gaussian prior: $\beta_{j}\sim\mathcal{N}(\mu_{j},\sigma_{j}^{2})$. The mean for feature $f_{j}$ is given by averaging this score across a set of paraphrased prompts T{T}:
$\mu_{j}=\frac{1}{|T|}\sum_{i\in T}\text{logit}_{j}(p_{i})$
This value reflects the LLM’s normalized preference for a positive versus negative association with the feature. Figure 1 demonstrates how LoID extracts informative priors from model logits. To capture uncertainty in the LLM’s belief, we define the standard deviation $\sigma_{j}$ as:
$\sigma_{j}^{2}=\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
Here, $logit_{i}(p_{j})$ denotes the logit preference score for the $i$th prompt and feature $j$. This design assigns higher variance to features for which the LLM shows inconsistent preferences across paraphrases, capturing greater uncertainty about their influence. We evaluated three variance computation methods and concluded that taking the standard deviation of logit preference scores as described above performs best (Section A.2.3).
@@ 3.4 Bayesian inference with LoID priors
With LoID priors, posterior inference is performed using a Laplace approximation, which approximates the posterior with a Gaussian centered at the maximum a posteriori (MAP) estimate:
$\hat{\beta}=\arg\max_{\beta}\log p(\beta\mid D_{\text{train}})$ (1)
The posterior is then locally approximated via a second-order Taylor expansion of the log-posterior:
$p(\beta\mid D_{\text{train}})\approx\mathcal{N}(\hat{\beta},\Sigma),\quad\Sigma=\left[\nabla^{2}_{\beta}\left(-\log p(\beta\mid D_{\text{train}})\right)\big|_{\beta=\hat{\beta}}\right]^{-1}$ (2)
This provides efficient uncertainty estimates without sampling, with LoID priors affecting both the MAP location and posterior curvature. All results are computed under this Gaussian approximation.
@@ 3.5 Evaluation
To evaluate LoID, we construct synthetic OOD splits from real-world datasets. We begin by creating a standard 80–20 random split, where the training portion serves as the in-distribution (IID) training set and the held-out portion serves as the IID test set. To simulate distribution shift, we derive an OOD training set by selecting a subset of the IID training data and systematically excluding specific categories of categorical features or ranges of numerical features. Section A.5 explains the synthetic distribution shift for each dataset in detail.
We verify the presence of distribution shift by comparing the predictive performance of a model trained on the IID training set with that of a model trained on the OOD training set, both evaluated on the same IID test set. A substantial degradation in performance when training on the OOD split indicates that the training data no longer adequately represents the true data distribution.
Given this setup, we then train Bayesian models with LoID priors using the OOD training set and evaluate their performance on the IID test set. Our primary objective is to assess the extent to which LoID priors can mitigate the performance gap introduced by distribution shift.
@@ 4.1 Datasets
Head and Neck Cancer Recurrence dataset
We compiled a binary classification dataset for Head and Neck Cancer Recurrence. This dataset includes comprehensive features about patient demographics, tumor characteristics, treatment regimens, and clinical staging, and predicts the likelihood of cancer recurrence following radiotherapy. The dataset encompasses 508 patients with head and neck cancers, capturing clinical variables including comorbidity indices (Charlson score), smoking history (pack-years), HPV status, primary tumor site localization, and detailed radiation therapy parameters such as treatment duration, prescription dose, and the critical time interval between biopsy and treatment initiation.
This dataset presents a significant class imbalance (390 non-recurrence vs. 118 recurrence cases), reflecting real-world clinical distributions. The dataset is particularly valuable for evaluating model performance under distributional shift, as we created an out-of-distribution test set focusing on elderly patients ($age\geq 75$). This datasets represents a clinically meaningful population that is not present in any LLM pre-training corpora. [^1] We will release this dataset upon publication.
Public datasets
We evaluate LoID on 14 publicly available tabular datasets for binary classification and regression: Adult Income Becker and Kohavi (1996), Cervical Cancer Fernandes et al. (2017), Chronic Kidney Disease Rubini et al. (2015), Bank Marketing Moro et al. (2014) Blood Transfusion Service Center Yeh (2008), Stroke Fedesoriano (2020), Heart Failure UCI Machine Learning Repository (2020), Diabetes Prediction Mustafatz (2023), Give Me Some Credit Kaggle (2011), Airfoil Self-Noise, Bike Sharing Fanaee-T (2013), Combined Cycle Power Plant Tfekci and Kaya (2014), Concrete Compressive Strength, Yeh (1998), and Communities and Crime Redmond (2002).
Preprocessing
@@
Across all datasets, we apply the same preprocessing pipeline: categorical features are one-hot encoded, missing categorical values are imputed with mode and missing numerical values are imputed with mean. For datasets with an excessive number of features—namely Bank Marketing, Breast Cancer Diagnostic, Cervical Cancer, Chronic Kidney Disease, Communities and Crime, Heart Failure, Haptitis, and Head and Neck Cancer Recurrence—we perform feature selection using Claude-3.5-Sonnet by providing feature descriptions and prompting it to return a list of the 15 most relevant features for the prediction task. The same reduced feature set is used consistently across all experimental conditions to ensure comparability. We apply z-score normalization to the features for AutoElicit and LoID, but not for LLMProcesses, following the authors’ recommendations.
Development datasets
@@
We designate the Stroke, Blood Donation, Heart Failure, and Adult Income datasets as development datasets for choosing an LLM and hyperparameter tuning A.2. The remaining 11 datasets are reserved for testing the generalization performance of the final LoID method.
4.2 Creating distribution shifts
For each dataset, we construct synthetic OOD splits as described in Section A.5. As shown in Table 1 in the P_IID (in-distribution train set performance), and P_OOD (out-of-distribution train set performance) columns, the model trained on the OOD train set noticeably underperforms compared to the model trained on the IID train set. This confirms that generalization is impaired and the OOD train set does not represent the full distribution of the dataset as well as the IID train set.
Dataset P_OOD P_IID N(0,1) LoID LLMProc AutoEl Gap (%)
Cap our method Requeima et al. (2024) Capstick et al. (2025)
Classification (AUC)
Head & Neck Cancer 0.542 0.998 0.567 0.696 0.540 0.609 33.8
Adult Income 0.414 0.809 0.51 0.745 0.603 0.510 83.8
Cervical Cancer 0.440 0.870 0.448 0.681 0.658 0.570 56.0
Kidney Disease 0.822 0.993 0.715 0.901 0.590 0.320 46.2
Diabetes 0.883 0.963 0.853 0.906 0.502 0.864 28.8
Stroke 0.608 0.842 0.618 0.694 0.500 0.670 36.5
Heart Failure 0.746 0.849 0.742 0.753 0.770 0.748 6.8
Bank Marketing 0.780 0.878 0.780 0.652 0.498 0.692 -130.6
Blood Donation 0.572 0.748 0.572 0.663 0.483 0.556 51.7
Good Credit Score 0.710 0.790 0.710 0.760 0.700 0.710 62.5
Regression (MSE)
Concrete Strength 469.20 98.14 165.72 120.59 136.57 1974.11 93.9
Communities & Crime 0.17 0.01 0.073 0.07 0.15 0.30 62.5
Airfoil Noise 77.16 23.69 51.11 60.51 47.80 62.21 68.9
Cycle Power 962.94 21.24 23.58 509.92 462.89 47.14 0.1
Bike Sharing 77858 19030 71890 78351 75650 71405 -1.8
Table 1: Performance comparison across 15 datasets. Gap Closed represents the percentage of the OOD performance gap recovered by our method, LoID. The percentage closed in calculated using: $\frac{|LoID-P\_OOD|}{|P\_IID-P\_OOD|}\times 100$